# Better Estimation of the KL Divergence Between Language Models

### v2

@@ E.2 Proof of Lemma 4
In the proof below, we consider the case where $M=1$. The proof easily generalizes to $M>1$ using the i.i.d. assumption. Additionally, we write $\delta_{\textsc{rb}}^{N}(\overline{\boldsymbol{Y}})$, rather than suppressing the argument to the estimator as we do in the main text. We begin by proving the unbiasedness property of the estimator.
$\operatorname*{\mathbb{E}}\big{[}\delta_{\textsc{rb}}^{N}\big{]}-\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}^{N}]$ $=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\Big{[}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\big{[}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\mid\overline{\boldsymbol{Y}}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\big{]}\Big{]}-\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}^{N}]$ (definition of $\delta_{\textsc{rb}}^{N}$) (38a)
$=\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\Big{[}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\big{[}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\mid\overline{\boldsymbol{Y}}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\big{]}\Big{]}-\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}^{N}]$ (linearity of expectation) (38b)
$=\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\right]-\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}^{N}]$ (law of total expectation) (38c)
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\right]-\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}^{N}]$ (linearity of expectation) (38d)
$=0.$ ($\delta_{\textsc{mc}}^{N}=\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}$) (38e)
Next, we prove the variance reduction property:
$\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{rb}}^{N}}\right]$ $=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left(\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\,\middle|\,\overline{\boldsymbol{Y}}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\right]-\operatorname*{\mathbb{E}}[\delta_{\textsc{rb}}^{N}]\right)^{2}\right]$ (39a)
(defintion of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(defintion of $\delta_{\textsc{rb}}^{N}$)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left(\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{n}}\left[\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{n})\,\middle|\,\overline{\boldsymbol{Y}}^{n}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\right]-\operatorname*{\mathbb{E}}[\delta_{\textsc{rb}}^{N}]\right)^{2}\right]$ (39b)
[…]
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left(\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{n}}\left[\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{n})-\operatorname*{\mathbb{E}}[\delta_{\textsc{rb}}^{N}]\,\middle|\,\overline{\boldsymbol{Y}}^{n}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\right]\right)^{2}\right]$ (39c)
(linearity of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of expectation)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left(\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{1}}\left[\cdots\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{N}}\left[\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{n})-\operatorname*{\mathbb{E}}[\delta_{\textsc{rb}}^{N}]\,\middle|\,\overline{\boldsymbol{Y}}^{N}_{<N}=\overline{\boldsymbol{Y}}^{\prime}_{<N}\right]\cdots\,\middle|\,\overline{\boldsymbol{Y}}^{1}_{<1}=\overline{\boldsymbol{Y}}^{\prime}_{<1}\right]\right)^{2}\right]$ (39d)
(linearity of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of expectation)}}\end{subarray}$
$\leq\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{1}}\left[\cdots\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{N}}\left[\left(\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{n})-\operatorname*{\mathbb{E}}[\delta_{\textsc{rb}}^{N}]\right)^{2}\,\middle|\,\overline{\boldsymbol{Y}}^{N}_{<N}=\overline{\boldsymbol{Y}}^{\prime}_{<N}\right]\cdots\,\middle|\,\overline{\boldsymbol{Y}}^{1}_{<1}=\overline{\boldsymbol{Y}}^{\prime}_{<1}\right]\right]$ (39e)
(Jensen’s $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Jensen's inequality)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{1}}\left[\cdots\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{N}}\left[\left(\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{n})-\operatorname*{\mathbb{E}}[\delta_{\textsc{rb}}^{N}]\right)^{2}\right]\right]$ (39f)
(law of total $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(law of total expectation)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\left(\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})-\operatorname*{\mathbb{E}}[\delta_{\textsc{rb}}^{N}]\right)^{2}\right]$ (39g)
[…]
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\left(\delta_{\textsc{mc}}^{N}(\overline{\boldsymbol{Y}})-\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}^{N}]\right)^{2}\right]$ (39h)
[…]
$=\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{mc}}^{N}}\right].$ (39i)
(definition of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(definition of variance)}}\end{subarray}$
E.3 Proof of Thm. 5
See 5
Proof.
First, in the following lemma, we prove that $\delta_{\textsc{mc}}$ can be decomposed into steps for the unbounded case as well. In the rest of the proof, we consider the special case of $M=1$. The proof easily generalizes to $M>1$ using the i.i.d. assumption. Additionally, we write $\delta_{\textsc{rb}}(\overline{\boldsymbol{Y}})$, rather than suppressing the argument to the estimator as we do in the main text. We begin with proving the unbiasedness of the estimator.
$\operatorname*{\mathbb{E}}\big{[}\delta_{\textsc{rb}}(\overline{\boldsymbol{Y}})\big{]}-\textsc{kl}(p\mid\mid q)$ $=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\Big{[}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\big{[}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\mid\overline{\boldsymbol{Y}}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\big{]}\Big{]}-\textsc{kl}(p\mid\mid q)$ (40a)
$=\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\Big{[}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\big{[}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\mid\overline{\boldsymbol{Y}}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\big{]}\Big{]}-\textsc{kl}(p\mid\mid q)$ (40b)
(Tonelli’s $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Tonelli's Theorem)}}\end{subarray}$
$=\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\right]-\textsc{kl}(p\mid\mid q)$ (40c)
[…]
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\right]-\textsc{kl}(p\mid\mid q)$ (40d)
(Fubini’s $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Fubini's Theorem,$\operatorname*{\mathbb{E}}\left[\sum_{n=1}^{\infty}|\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})|\right]<\infty$)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\delta_{\textsc{mc}}(\overline{\boldsymbol{Y}})\right]-\textsc{kl}(p\mid\mid q)$ (40e)
[…]
$=0.$ (40f)
(unbiasedness of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(unbiasedness of $\delta_{\textsc{mc}}$)}}\end{subarray}$
Finally, we prove the variance reduction property.
$\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{rb}}}\right]$ $=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left(\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\big{[}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\mid\overline{\boldsymbol{Y}}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\big{]}-\textsc{kl}(p\mid\mid q)\right)^{2}\right]$ (41a)
(Definition of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Definition of $\delta_{\textsc{rb}}(\boldsymbol{Y})$)}}\end{subarray}$
$=\lim_{N\rightarrow\infty}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left(\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\big{[}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})\mid\overline{\boldsymbol{Y}}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\big{]}-\textsc{kl}(p\mid\mid q)\right)^{2}\right]$ (41b)
(dominated convergence theorem, $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(dominated convergence theorem, $\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{rb}}}\right]<\infty$)}}\end{subarray}$
$\leq\lim_{N\rightarrow\infty}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\left(\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})-\textsc{kl}(p,q)\right)^{2}\right]$ (41c)
[…]
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\lim_{N\rightarrow\infty}\left(\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}})-\textsc{kl}(p,q)\right)^{2}\right]$ (41d)
(dominated convergence theorem, $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(dominated convergence theorem, $\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{mc}}}\right]<\infty$)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\left(\delta_{\textsc{mc}}(\overline{\boldsymbol{Y}})-\textsc{kl}(p\mid\mid q)\right)^{2}\right]$ (41e)
[…]
$=\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{mc}}}\right].$ (41f)
Appendix F Rao–Blackwellized Estimator of the Gradient
See 7
Proof.
@@ Appendix F Rao–Blackwellized Estimator of the Gradient
$\nabla_{\boldsymbol{\theta}}\textsc{kl}(p\mid\mid q)$ $=\sum_{\boldsymbol{y}\in\Sigma^{*}}\textsc{kl}\left(\vec{p}_{\boldsymbol{\theta}}(\cdot\mid\boldsymbol{y})\mid\mid\vec{q}(\cdot\mid\boldsymbol{y})\right)\nabla_{\boldsymbol{\theta}}\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})+\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\nabla_{\boldsymbol{\theta}}\textsc{kl}\left(\vec{p}_{\boldsymbol{\theta}}(\cdot\mid\boldsymbol{y})\mid\mid\vec{q}(\cdot\mid\boldsymbol{y})\right)$ (42a)
@@
F.1 Variance Reduction Proof
The proof structure is as follows: We first prove that the inequality holds when we constrain $\overline{\boldsymbol{Y}}$ to have length less than or equal to $N$. We then generalize to the infinite-length sequences by analyzing as $N\rightarrow\infty$. We begin with defining the length-conditioned MC and RB estimators. Let $\nabla\delta_{\textsc{mc}}^{N}$ be the length-conditioned MC estimator of the gradient:
$\nabla\delta_{\textsc{mc}}^{N}=\sum_{n=1}^{N}\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}_{<n}^{(m)})}{\vec{q}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}_{<n}^{(m)})}\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq N}^{(m)}).$ (43)
Let $\nabla\delta_{\textsc{rb}}^{N}$ be the length-conditioned RB estimator:
$\nabla\delta_{\textsc{rb}}^{N}=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}^{(m)}_{<n})\right)\right].$ (44)
@@ F.1 Variance Reduction Proof
The length-conditioned MC estimator of the gradient $\nabla\delta_{\textsc{mc}}^{N}$ converges to $\nabla\delta_{\textsc{mc}}$ as $N$ goes to $\infty$, i.e., $\lim_{N\rightarrow\infty}\nabla\delta_{\textsc{mc}}^{N}=\nabla\delta_{\textsc{mc}}$.
@@
$\nabla\delta_{\textsc{mc}}$ $=\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{\vec{q}(\boldsymbol{Y}^{(m)})}\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})$ (45)
@@
$=\lim_{N\rightarrow\infty}\nabla\delta_{\textsc{mc}}^{N}.$ (51)
@@
$\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n})=\frac{\nabla_{\boldsymbol{\theta}}\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n})}{\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n})}$ $=\sum_{\overline{\boldsymbol{y}}^{\prime}\in\overline{\Sigma}^{N-n}}\frac{\nabla_{\boldsymbol{\theta}}\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n}\overline{\boldsymbol{y}}^{\prime})}{\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n})}$ (53a)
$=\sum_{\overline{\boldsymbol{y}}^{\prime}\in\overline{\Sigma}^{N-n}}\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n}\overline{\boldsymbol{y}}^{\prime})}{\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n})}\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n}\overline{\boldsymbol{y}}^{\prime})$ (53b)
$=\sum_{\overline{\boldsymbol{y}}^{\prime}\in\overline{\Sigma}^{N-n}}\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{y}}^{\prime}\mid\overline{\boldsymbol{Y}}_{\leq n})\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n})}{\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n})}\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n}\overline{\boldsymbol{y}}^{\prime})$ (53c)
$=\sum_{\overline{\boldsymbol{y}}^{\prime}\in\overline{\Sigma}^{N-n}}\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{y}}^{\prime}\mid\overline{\boldsymbol{Y}}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n}\overline{\boldsymbol{y}}^{\prime})$ (53d)
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq N}^{\prime})\,\middle|\,\overline{\boldsymbol{Y}}^{\prime}_{\leq n}=\overline{\boldsymbol{Y}}_{\leq n}\right]$ (53e)
@@
Let $\nabla\delta_{\textsc{mc}}^{N}$ be the length-conditioned MC estimator of the gradient defined in Eq. 43 and $\nabla\delta_{\textsc{rb}}^{N}$ the length-conditioned RB estimator of the gradient defined in Eq. 44. Define $G_{N}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\operatorname*{\mathbb{E}}[\nabla\delta_{\textsc{mc}}^{N}]=\operatorname*{\mathbb{E}}[\nabla\delta_{\textsc{rb}}^{N}]$. We have
$\operatorname*{\mathbb{E}}\left[\left\lVert\nabla\delta_{\textsc{rb}}^{N}-G_{N}\right\rVert^{2}\right]\leq\operatorname*{\mathbb{E}}\left[\left\lVert\nabla\delta_{\textsc{mc}}^{N}-G_{N}\right\rVert^{2}\right].$ (54)

### v3

@@ Appendix D Rao–Blackwellized Estimator of the Gradient
$\nabla_{\boldsymbol{\theta}}\textsc{kl}(p\mid\mid q)$
@@
$=\sum_{\boldsymbol{y}\in\Sigma^{*}}\textsc{kl}\left(\vec{p}_{\boldsymbol{\theta}}(\cdot\mid\boldsymbol{y})\mid\mid\vec{q}(\cdot\mid\boldsymbol{y})\right)\nabla_{\boldsymbol{\theta}}\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})+\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\nabla_{\boldsymbol{\theta}}\textsc{kl}\left(\vec{p}_{\boldsymbol{\theta}}(\cdot\mid\boldsymbol{y})\mid\mid\vec{q}(\cdot\mid\boldsymbol{y})\right)$ (40a)
@@
D.1 Variance-Reduction Proof
The proof structure is as follows: We first prove that the inequality holds when we constrain $\overline{\boldsymbol{Y}}$} to have length less than or equal to NN. We then generalize to the infinite-length sequences by analyzing as $N\rightarrow\infty$. We begin with defining the truncated MC and RB estimators. Let $\boldsymbol{\delta}_{\textsc{mc}}^{(N)}$ be the truncated MC estimator of the gradient:
$\boldsymbol{\delta}_{\textsc{mc}}^{(N)}=\sum_{n=1}^{N}\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}_{\boldsymbol{\theta}}({\overline{Y}}_{n}^{(m)}\mid{\overline{\boldsymbol{Y}}}_{<n}^{(m)})}{\vec{q}({\overline{Y}}_{n}^{(m)}\mid{\overline{\boldsymbol{Y}}}_{<n}^{(m)})}\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq N}^{(m)}).$ (41)
Let $\boldsymbol{\delta}_{\textsc{rb}}^{(N)}$ be the truncated RB estimator:
$\boldsymbol{\delta}_{\textsc{rb}}^{(N)}=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{{\overline{Y}}_{n}}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}({\overline{Y}}_{n}\mid{\overline{\boldsymbol{Y}}}^{(m)}_{<n})}{\vec{q}({\overline{Y}}_{n}\mid{\overline{\boldsymbol{Y}}}^{(m)}_{<n})}\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{Y}}_{n}\mid{\overline{\boldsymbol{Y}}}^{(m)}_{<n})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}^{(m)}_{<n})\right)\right$]. (42)
@@ D.1 Variance-Reduction Proof
The truncated MC estimator of the gradient $\boldsymbol{\delta}_{\textsc{mc}}^{(N)}$ converges to $\boldsymbol{\delta}_{\textsc{mc}}$ as NN goes to $\infty$, i.e., $\lim_{N\rightarrow\infty}\boldsymbol{\delta}_{\textsc{mc}}^{(N)}=\boldsymbol{\delta}_{\textsc{mc}}$.
@@
$\boldsymbol{\delta}_{\textsc{mc}}$ $=\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{\vec{q}(\boldsymbol{Y}^{(m)})}\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})$ (43)
@@
$=\lim_{N\rightarrow\infty}\boldsymbol{\delta}_{\textsc{mc}}^{(N)}.$ (49)
@@
$\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n})$ $=\frac{\nabla_{\boldsymbol{\theta}}\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n})}{\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n})}$ (51a)
$=\sum_{{\overline{\boldsymbol{y}}}^{\prime}\in\overline{\Sigma}^{N-n}}\frac{\nabla_{\boldsymbol{\theta}}\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n}{\overline{\boldsymbol{y}}}^{\prime})}{\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n})}$ (51b)
$=\sum_{{\overline{\boldsymbol{y}}}^{\prime}\in\overline{\Sigma}^{N-n}}\frac{\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n}{\overline{\boldsymbol{y}}}^{\prime})}{\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n})}\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n}{\overline{\boldsymbol{y}}}^{\prime})$ (51c)
$=\sum_{{\overline{\boldsymbol{y}}}^{\prime}\in\overline{\Sigma}^{N-n}}\frac{\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{y}}}^{\prime}\mid{\overline{\boldsymbol{Y}}}_{\leq n})\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n})}{\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n})}\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n}{\overline{\boldsymbol{y}}}^{\prime})$ (51d)
$=\sum_{{\overline{\boldsymbol{y}}}^{\prime}\in\overline{\Sigma}^{N-n}}\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{y}}}^{\prime}\mid{\overline{\boldsymbol{Y}}}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n}{\overline{\boldsymbol{y}}}^{\prime})$ (51e)
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime}}\left[\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq N}^{\prime})\,\middle|\,{\overline{\boldsymbol{Y}}}^{\prime}_{\leq n}={\overline{\boldsymbol{Y}}}_{\leq n}\right]$ (51f)
@@
Let $\boldsymbol{\delta}_{\textsc{mc}}^{(N)}$ be the truncated MC estimator of the gradient defined in Eq. 41 and $\boldsymbol{\delta}_{\textsc{rb}}^{(N)}$ the truncated RB estimator of the gradient defined in Eq. 42. Define $\boldsymbol{G}^{N}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\operatorname*{\mathbb{E}}[\boldsymbol{\delta}_{\textsc{mc}}^{(N)}]=\operatorname*{\mathbb{E}}[\boldsymbol{\delta}_{\textsc{rb}}^{(N)}]$. We have
$\operatorname*{\mathbb{E}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{rb}}^{(N)}-\boldsymbol{G}^{N}\right\rVert^{2}\right]\leq\operatorname*{\mathbb{E}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{mc}}^{(N)}-\boldsymbol{G}^{N}\right\rVert^{2}\right].$ (52)